in a group of 20 adults, 4 out of 7 women and 2 out of the 13 men wear glasses. What is the probability that a person chosen at random from the group is a woman or someone who wear glasses? 


1
Expert's answer
2021-05-07T09:16:02-0400

There are 7 women in the group and 13 men. 6 people wear glasses whereas 13 does not. There are 4 women who wear glasses.

Let A be the event that one is a woman and B be the event that they wear glasses.

probability that one is a female or wears glasses"P(A\cup B)"

"P(A \cup B)= P(A)+P(B)-P(A \cap B)"

"=\frac 7 {20}+\frac 6 {20}-\frac 4 {20}"

"=\frac 9 {20}"